. Answer the following questions about the Ni2+ atom.

(a) What is the electron configuration?

(b) What are the possible term symbols for a Ni2+ atom for the configuration in part (a)?

(c) Of the possible term symbols in part (b), which would you predict to be the ground state? Why?
Answer #1

(1) Ni2+ : 1s2 2s2 2p6 3s2 3p6 3d8 or

you can shorten it to [Ar]3d8

No 2+ has two electrons less than Ni ( that is why Ni2+ is positively charged).

So when writing electron configuration for Ni2+we exclude last two electrones from the last shell of Ni electron confgiruation.

(2) To find the ground state term symbol, we should use symmetry and group theory arguments.

The electron configuration of Ni²⁺ is - [Ar]3d⁸.

ml = ↑↓ +2 ↑↓+1 ↑↓0 ↑-1 ↑-2

So L = 0 + 0 + 2 + 2 + 1 + 1 + 0 + 0 + –1 + –2 = 3 (F state),

and S = ½ + –½ + ½ + –½ + ½ + –½ + ½ + –½ + ½ + ½ = 1 (2S + 1 = 3).

This gives a term symbol of ³F.

Now if the confugration become [Ar] 3d7 4s1 then state symbol become L will remain same and S will become 2 then 2s+1 will be 5 so the term symbol is 5F.

similarly isf confugration become [Ar] 3d6 4s2 then  

L= 2 (D state)

And symbol will be 5D

(3) among above the ground state symbol 3F will be most stable because in this the value of energy means (2S +1) is minimum. Which leads the stable confugration.